# FuseMoE: Mixture-of-Experts Transformers for Fleximodal Fusion

### v2

@@ D.2 PAM Dataset
We follow the preprocessing procedure from [98] as published from their official GitHub repository [^1] https://github.com/mims-harvard/Raindrop/tree/main/PAMdata.
@@ E.1 Unified Temporal Discretization Embeddings
Unlike the embeddings in chest X-rays, clinical notes, and ECGs, vitals/lab/time-series values present temporal irregularity across dimensions. That is, for the former three modalities, each dimension of the corresponding is observed at each irregular time point $\bm{\tau}$. By contrast, the sampling for vitals/labs is irregular in both within and across dimensions. For example, we might observe heart rate values sampled at times $\bm{\tau}_{\text{HR}}=\{0,0.2,0.8,1.2,2.8\}$ and glucose values sampled at time $\bm{\tau}_{\text{Glu}}=\{0.1,0.7,3.4\}$. Given this unique challenge present in vitals/labs, we adapt the Unified Temporal Discretization Embedding (UTDE) approach described in [99], which combines the mTAND approach described in Section 2.2 with a simpler imputation-based discretization scheme. Specifically, given a set of $t$ observations $\mathbf{x}\in\mathbb{R}^{t}$ observed at irregular times $\bm{\tau}\in\mathbb{R}^{t}$, we a simple imputation scheme to discretize $\mathbf{x}$ into target bins $\bm{\gamma}$ (e.g., $\bm{\gamma}=\{0,1,2,...,\gamma\}$. Specifically, given bin value $\gamma_{i}\in\bm{\gamma}$, we apply the following rules:
@@ E.2 Unifying imputation and mTAND embeddings
We combined simple imputation and mTAND embeddings via a gating function $\mathbf{g}$. Following [99], we let $\mathbf{E}_{\text{mTAND}}\in\mathbb{R}^{\gamma\times d_{h}}$ denote the mTAND embeddings for vitals/labs derived from the process described in Section 2.2 and let $\mathbf{E}_{\text{Imp}}\in\mathbb{R}^{\gamma\times d_{h}}$ denote the simple imputations from the process described above. We use each of these discretization embeddings to derive a final set of embeddings for vitals/labs $\mathbf{E}_{\text{vitals}}$ via a one-layer MLP gating function $f$. Specifically, we let $\mathbf{g}=f(\mathbf{E}_{\text{Imp}}\oplus\mathbf{E}_{\text{mTAND}})$, where $\oplus$ denotes the concatenation operator. We then calculate $\mathbf{E}_{\text{vitals}}$ as
@@ F.1 MISTS
This approach, from [99], casts time series and clinical notes as multivariate, irregularly-sampled time series (MISTS) and uses layers of self- and cross-attention to fuse modalities. The method uses a Time2Vec [43] encoding scheme to represent the irregularity of observation times. We use the same hyperparameters as in the original paper (e.g., $3$ self- and cross-attention blocks, $128$-dimensional time embedding, etc.).
@@ F.2 MulT
This model from [88] relies on multiple stacks of pairwise and bidirectional cross-modal attention blocks (without a self-attention mechanism) to attend to low-level features. The results of cross-modal attention are then sent to modality-specific transformers, concatenated, and used to make predictions.
@@ F.3 MAG
This method introduces the Multimodal Adaptation Gate (MAG) as an extension to BERT and XLNet, allowing these pre-trained models to incorporate visual and acoustic data during fine-tuning. By generating a modality-conditioned shift in their internal representations, MAG enables enhanced sentiment analysis performance on multimodal datasets, achieving human-level accuracy in the field [72].
@@ F.4 TFN
The proposed Tensor Fusion Network approach (TFN) integrates three core components: Modality Embedding Subnetworks for generating rich embeddings from unimodal inputs, a Tensor Fusion Layer for capturing all levels of modality interactions through a 3-fold Cartesian product, and a Sentiment Inference Subnetwork tailored to perform sentiment analysis based on the fusion layer’s output [94].
@@ F.5 HAIM
The multimodal fusion approach detailed by [81] extracts a single set of features for each ICU stay, and uses this to predict the outcome of interest (in-hospital mortality, etc.). For vitals/lab values, the authors extract a set of 11 generic time series features: signal length, maximum, minimum, mean, median, SD, variance, number of peaks, and average time-series slope and piece-wise change over time of these metrics. This is done independently for each of the 30 events, leading to $30\times 11=330$ vital/lab features per ICU stay. To provide a fair comparison with our method, we only include the most recent five notes and $128$ vitals measurements in calculating embeddings. We only include entries for which all modalities are observed. For note/X-ray/ECG embeddings, we compute the mean embedding across all observations occurring during the specified time frame (i.e., the first 48 hours of 48-IHM and LOS, the entire stay for PHE). As with our method, we standardize scale values based on the training set. [81] uses an XGBoost [12] classifier to predict the outcomes of interest. We follow the hyperparameter optimization approach described in the paper. Specifically, we conduct a grid search across the following sets of hyperparameters: max depth =$=\{5,6,7,8\}$, number of estimators =$=\{200,300\}$, learning rate =$=\{0.3,0.1,0.05\}$. Hyperparameters are selected based on the maximum AU-ROC from five-fold cross-validation.
@@ F.6 Implementation
We integrate F.1 through F.4 into our workflow using the implementation provided by [99]. For F.5, we adapt the time series (e.g., series variance, mean, etc.) feature extraction and model fitting code from the repository released by the corresponding paper. The original paper doesn’t use ECG waveforms, so we adopt a similar approach to ECG embeddings as with image and note embeddings, and take the mean value of the latent vector across all included observations.
@@ G.2 Hyper-Parameters
All parameters can be found in Table 7 and the codebase submitted in the zip file.
@@ H.1 FlexiModal Experiments
Table 8: Comparison of FuseMoE-based methods (gray) and baselines, utilizing vital signs and clinical notes of MIMIC-III. The best results are highlighted in bold font, and the second-best results are underlined. All results are averaged across 5 random runs. Since HAIM is not designed for the MIMIC-III dataset, we use the concatenation method from e.g. [45] as a replacement.
@@ H.2 Ablation Study on FuseMoE Building Blocks
    utilizing solely the mTAND module [80] within the time-series irregularity encoder
@@
    implementing the SeFT method [31] as an irregularity encoder
@@
    adopting the RAINDROP method [98] as an irregularity encoder
@@
In Figure 8, we evaluate the impact of various time-series encoders on the performance of the FuseMoE framework. The original FuseMoE framework feeds time-series embeddings obtained from the irregularity encoder into the Transformer [90] and extracts the last hidden states of the Transformer output to pass through fully connected layers to make predictions. Our baseline approaches include CNN [47] and LSTM [30] to encode time-series embeddings from the irregularity encoder.
@@
In Figure 9, we assess the effect of different CXR encoders on the FuseMoE framework. Currently, the FuseMoE framework incorporates DenseNet-121 as the feature extractor for CXR images before their integration into the mTAND module. This setup is compared with the application of the state-of-the-art vision transformer (ViT-B) [17] as an alternative CXR encoder.
@@
In Figure 10, we evaluate the influence of text encoders on the FuseMoE framework. Currently, FuseMoE incorporates Clinical-Longformer [49] as the text encoder before integrating it into the mTAND module. This setup is compared with other state-of-the-art text encoders: GRU-D [10], FT-LSTM [97], and HierTrans [67].
@@ H.3 Ablation Study on MoE Architecture
We then conducted ablation studies to explore the efficiency and effectiveness of MoE architecture on model performance. We mainly use MIMIC-IV as our test bed. Figure 12(a) examines the computational efficiency and resource utilization, positioning FuseMoE approximately in the middle of the comparison. Despite the increase in model parameters due to the incorporation of the MoE layer, its sparse nature does not significantly escalate the computational load. Figure 12(b) illustrates the correlation between the number of experts and task performance across different modalities. Generally, performance improves with the addition of more experts, plateauing once the count exceeds 16. To achieve a compromise between performance and computational expense, we opted to utilize the top 4 experts out of 16 in our experiments. Figure 12(c) studies the influence of each modality on the top-$k$ chosen experts. For every expert selected, we calculate the number of samples that include a specific modality, weighted by corresponding weight factors from the gating functions. The outcomes are subsequently normalized across modalities. The analysis reveals that predictions across all tasks heavily depend on vital signs and clinical notes. This reliance is attributed to the abundant samples in these two modalities. Despite the notably smaller quantity of CXR, they play more significant roles in the 25-PHE and 48-IHM tasks, which aligns with our findings in Table 5.
@@ Appendix I Details on Numerical Experiments
We conduct multiple numerical experiments to illustrate the theoretical convergence rates of the MLE $\widehat{G}_{n}$ to the true mixing measure $G_{*}$ under both the exact-specified and the over-specified settings.
@@ I.1 Experimental Setup
Then, we generate i.i.d samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$ by first sampling $X_{i}$’s from the uniform distribution $\mathrm{Uniform}[0,1]$ and then sampling $Y_{i}$’s from the true conditional density $g_{G_{*}}(Y|X)$ of the Laplace gating Gaussian mixture of experts (MoE) given in equation 4. Throughout the following experiments, we will consider only the scenario when $K=1$, that is, we choose the best expert from the true conditional density $g_{G_{*}}(Y|X)$.
Maximum Likelihood Estimation (MLE). A popular approach to determining the MLE $\widehat{G}_{n}$ for each set of samples is to use the EM algorithm [15]. However, since there are not any closed-form expressions for updating the gating parameters $\beta_{0i},\beta_{1i}$ in the maximization steps, we have to leverage an EM-based numerical scheme, which was previously used in [63]. In particular, we utilize a simple coordinate gradient descent algorithm in the maximization steps. Additionally, we select the convergence criterion of $\epsilon=10^{-6}$ and run a maximum of 2000 EM iterations.
@@ I.2 Exact-specified Setting
Under the exact-specified settings, we conduct 5 sample generations for each configuration, across a spectrum of 10 different sample sizes $n$ ranging from $10^{3}$ to $10^{5}$. It can be seen from Figure 3 (left) that the MLE $\widehat{G}_{n}$ empirically converges to the true mixing measure $G_{*}$ under the Voronoi metric $\mathcal{D}_{1}$ at the rate of order $\widetilde{\mathcal{O}}(n^{-0.49})$, which matches the theoretical parametric convergence rate established in Theorem J.2.
@@ I.3 Over-specified Setting
Under the over-specified settings, we continue to generate 5 samples of size $n$ for each setting, given 10 different choices of sample size $n\in[10^{3},10^{5}]$. From Figure 3 (right), we observe that the MLE $\widehat{G}_{n}$ empirically converges to $G_{*}$ under the Voronoi metric $\mathcal{D}_{2}$ at the rate of order $\widetilde{\mathcal{O}}(n^{-0.47})$, which aligns with the theoretical parametric convergence rate established in Theorem 3.2.
@@
In this appendix, we study the theoretical behaviors of the MLE under the exact-specified setting, i.e., $k=k_{*}$, of the Laplace gating Gaussian MoE. We demonstrate that under the exact-specified setting, the rate of estimated conditional density function $p_{\widehat{G}_{n}}$ to $p_{G_{*}}$ is parametric $O(n^{-1/2})$ (up to some logarithmic factor).
@@
Proof of Theorem J.1 is in Appendix K.1. The result of Theorem J.1 indicates that as long as we can establish the lower bound of the total variation distance between $p_{\widehat{G}_{n}}$ and $p_{G_{*}}$ based on certain loss function between the MLE $\widehat{G}_{n}$ and the true mixing measure $G_{*}$, we directly achieve the rate of the MLE under that loss function.
Voronoi Loss We now define that loss function between the MLE and the true mixing measure. Given some mixing measure $G:=\sum_{i=1}^{k}\exp(\beta_{i})\delta_{(W_{i},a_{i},b_{i},\nu_{i})}$, we distribute its components to the following Voronoi cells which are generated by the support of the true mixing measure $G_{*}:=\sum_{j=1}^{k_{*}}\exp(\beta_{j}^{*})\delta_{(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})}$:
$\mathcal{A}_{j}\equiv\mathcal{A}_{j}(G):=\{i\in[k]:\|\theta_{i}-\theta^{*}_{j}\|\leq\|\theta_{i}-\theta^{*}_{\ell}\|,\ \forall\ell\neq j\},$
where $\theta_{i}:=(W_{i},a_{i},b_{i},\nu_{i})$ and $\theta^{*}_{j}:=(W_{j}^{*},a_{j}^{*},b_{j}^{*},\nu_{j}^{*})$ for any $1\leq i\leq k$ and $1\leq j\leq k_{*}$. Based on these Voronoi cells, we propose the following Voronoi loss function for the exact-specified setting:
$\mathcal{D}_{1}(G,G_{*}):=\max\Bigg{[}\Big{|}\sum_{i\in\mathcal{A}_{\tau_{j}}}\exp(\beta_{i})-\exp(\beta^{*}_{\tau_{j}})\Big{|}+\sum_{j\in[K]:|\mathcal{A}_{j}|=1}\sum_{i\in\mathcal{A}_{\tau_{j}}}\exp(\beta_{i})\Phi_{i\tau_{j}}(1,1,1,1)\Bigg{]},$ (10)
where the maximum in the definition of Voronoi loss function $\mathcal{D}_{1}$ is for $\{\tau_{1},\tau_{2},\ldots,\tau_{K}\}\subseteq[k_{*}]$. Furthermore, for any $(\rho_{1},\rho_{2},\rho_{3},\rho_{4})\in\mathbb{R}^{4}$, we define $\Phi_{i\tau_{j}}(\rho_{1},\rho_{2},\rho_{3},\rho_{4})=\|W_{i}-W_{\tau_{j}}^{*}\|^{\rho_{1}}+\|a_{i}-a_{\tau_{j}}^{*}\|^{\rho_{2}}+|b_{i}-b_{\tau_{j}}^{*}|^{\rho_{3}}+|\nu_{i}-\nu_{\tau_{j}}^{*}|^{\rho_{4}}$ for any $i\in\mathcal{A}_{\tau_{j}}$ and $j\in[K]$. We demonstrate in the following theorem that the rate of MLE to the true mixing measure under the Voronoi loss function $\mathcal{D}_{1}$ is $\mathcal{O}(n^{-1/2})$ (up to some logarithmic factor).
@@
Proof of Theorem J.2 is in Appendix K.2. The convergence rate of MLE under the Voronoi loss function $\mathcal{D}_{1}$ implies that the rates of estimating the true parameters $W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*}$ are also $\mathcal{O}(n^{-1/2})$ (up to logarithmic factors). These rates are comparable to those under the exact-specified setting of softmax gating Gaussian MoE (cf. Theorem 1 in  [64]).

### v4

@@ D.2 PAM Dataset
We follow the preprocessing procedure from Zhang et al. [2021] as published from their official GitHub repository [^1] https://github.com/mims-harvard/Raindrop/tree/main/PAMdata.
@@ E.1 Unified Temporal Discretization Embeddings
Unlike the embeddings in chest X-rays, clinical notes, and ECGs, vitals/lab/time-series values present temporal irregularity across dimensions. That is, for the former three modalities, each dimension of the corresponding is observed at each irregular time point $\bm{\tau}$. By contrast, the sampling for vitals/labs is irregular in both within and across dimensions. For example, we might observe heart rate values sampled at times $\bm{\tau}_{\text{HR}}=\{0,0.2,0.8,1.2,2.8\}$ and glucose values sampled at time $\bm{\tau}_{\text{Glu}}=\{0.1,0.7,3.4\}$. Given this unique challenge present in vitals/labs, we adapt the Unified Temporal Discretization Embedding (UTDE) approach described in Zhang et al. [2023], which combines the mTAND approach described in Section 2.2 with a simpler imputation-based discretization scheme. Specifically, given a set of $t$ observations $\mathbf{x}\in\mathbb{R}^{t}$ observed at irregular times $\bm{\tau}\in\mathbb{R}^{t}$, we a simple imputation scheme to discretize $\mathbf{x}$ into target bins $\bm{\gamma}$ (e.g., $\bm{\gamma}=\{0,1,2,...,\gamma\}$. Specifically, given bin value $\gamma_{i}\in\bm{\gamma}$, we apply the following rules:
@@ E.2 Unifying imputation and mTAND embeddings
We combined simple imputation and mTAND embeddings via a gating function $\mathbf{g}$. Following Zhang et al. [2023], we let $\mathbf{E}_{\text{mTAND}}\in\mathbb{R}^{\gamma\times d_{h}}$ denote the mTAND embeddings for vitals/labs derived from the process described in Section 2.2 and let $\mathbf{E}_{\text{Imp}}\in\mathbb{R}^{\gamma\times d_{h}}$ denote the simple imputations from the process described above. We use each of these discretization embeddings to derive a final set of embeddings for vitals/labs $\mathbf{E}_{\text{vitals}}$ via a one-layer MLP gating function $f$. Specifically, we let $\mathbf{g}=f(\mathbf{E}_{\text{Imp}}\oplus\mathbf{E}_{\text{mTAND}})$, where $\oplus$ denotes the concatenation operator. We then calculate $\mathbf{E}_{\text{vitals}}$ as
@@ F.1 MISTS
This approach, from Zhang et al. [2023], casts time series and clinical notes as multivariate, irregularly-sampled time series (MISTS) and uses layers of self- and cross-attention to fuse modalities. The method uses a Time2Vec Kazemi et al. [2019] encoding scheme to represent the irregularity of observation times. We use the same hyperparameters as in the original paper (e.g., $3$ self- and cross-attention blocks, $128$-dimensional time embedding, etc.).
@@ F.2 MulT
This model from Tsai et al. [2019] relies on multiple stacks of pairwise and bidirectional cross-modal attention blocks (without a self-attention mechanism) to attend to low-level features. The results of cross-modal attention are then sent to modality-specific transformers, concatenated, and used to make predictions.
@@ F.3 MAG
This method introduces the Multimodal Adaptation Gate (MAG) as an extension to BERT and XLNet, allowing these pre-trained models to incorporate visual and acoustic data during fine-tuning. By generating a modality-conditioned shift in their internal representations, MAG enables enhanced sentiment analysis performance on multimodal datasets, achieving human-level accuracy in the field Rahman et al. [2020].
@@ F.4 TFN
The proposed Tensor Fusion Network approach (TFN) integrates three core components: Modality Embedding Subnetworks for generating rich embeddings from unimodal inputs, a Tensor Fusion Layer for capturing all levels of modality interactions through a 3-fold Cartesian product, and a Sentiment Inference Subnetwork tailored to perform sentiment analysis based on the fusion layer’s output Zadeh et al. [2017].
@@ F.5 HAIM
The multimodal fusion approach detailed by Soenksen et al. [2022] extracts a single set of features for each ICU stay, and uses this to predict the outcome of interest (in-hospital mortality, etc.). For vitals/lab values, the authors extract a set of 11 generic time series features: signal length, maximum, minimum, mean, median, SD, variance, number of peaks, and average time-series slope and piece-wise change over time of these metrics. This is done independently for each of the 30 events, leading to $30\times 11=330$ vital/lab features per ICU stay. To provide a fair comparison with our method, we only include the most recent five notes and $128$ vitals measurements in calculating embeddings. We only include entries for which all modalities are observed. For note/X-ray/ECG embeddings, we compute the mean embedding across all observations occurring during the specified time frame (i.e., the first 48 hours of 48-IHM and LOS, the entire stay for PHE). As with our method, we standardize scale values based on the training set. Soenksen et al. [2022] uses an XGBoost Chen and Guestrin [2016] classifier to predict the outcomes of interest. We follow the hyperparameter optimization approach described in the paper. Specifically, we conduct a grid search across the following sets of hyperparameters: max depth =$=\{5,6,7,8\}$, number of estimators =$=\{200,300\}$, learning rate =$=\{0.3,0.1,0.05\}$. Hyperparameters are selected based on the maximum AU-ROC from five-fold cross-validation.
@@ F.6 Implementation
We integrate F.1 through F.4 into our workflow using the implementation provided by Zhang et al. [2023]. For F.5, we adapt the time series (e.g., series variance, mean, etc.) feature extraction and model fitting code from the repository released by the corresponding paper. The original paper doesn’t use ECG waveforms, so we adopt a similar approach to ECG embeddings as with image and note embeddings, and take the mean value of the latent vector across all included observations.
@@ G.2 Hyper-Parameters
The set of parameters we used for experiments can be found in Table 8.
@@ H.1 FlexiModal Experiments
Table 9: Comparison of FuseMoE-based methods (gray) and baselines, utilizing vital signs and clinical notes of MIMIC-III. The best results are highlighted in bold font, and the second-best results are underlined. All results are averaged across 5 random runs. Since HAIM is not designed for the MIMIC-III dataset, we use the concatenation method from e.g. Khadanga et al. [2019] as a replacement.
@@ H.2 Ablation Study on FuseMoE Building Blocks
    utilizing solely the mTAND module [Shukla and Marlin, 2021] within the time-series irregularity encoder
@@
    implementing the SeFT method [Horn et al., 2020] as an irregularity encoder
@@
    adopting the RAINDROP method [Zhang et al., 2021] as an irregularity encoder
@@
In Figure 8, we evaluate the impact of various time-series encoders on the performance of the FuseMoE framework. The original FuseMoE framework feeds time-series embeddings obtained from the irregularity encoder into the Transformer [Vaswani et al., 2017] and extracts the last hidden states of the Transformer output to pass through fully connected layers to make predictions. Our baseline approaches include CNN [LeCun et al., 1998] and LSTM [Hochreiter and Schmidhuber, 1997] to encode time-series embeddings from the irregularity encoder.
@@
In Figure 9, we assess the effect of different CXR encoders on the FuseMoE framework. Currently, the FuseMoE framework incorporates DenseNet-121 as the feature extractor for CXR images before their integration into the mTAND module. This setup is compared with the application of the state-of-the-art vision transformer (ViT-B) [Dosovitskiy et al., 2020] as an alternative CXR encoder.
@@
In Figure 10, we evaluate the influence of text encoders on the FuseMoE framework. Currently, FuseMoE incorporates Clinical-Longformer [Li et al., 2022b] as the text encoder before integrating it into the mTAND module. This setup is compared with other state-of-the-art text encoders: GRU-D [Che et al., 2018], FT-LSTM [Zhang et al., 2020], and HierTrans [Pappagari et al., 2019].
@@ H.3 Ablation Study on MoE Architecture
Figure 13: Modality weight composition of CMU-MOSI and CMU-MOSEI datasets.
We then conducted ablation studies to explore the efficiency and effectiveness of MoE architecture on model performance. We mainly use MIMIC-IV as our test bed. Figure 12(a) examines the computational efficiency and resource utilization, positioning FuseMoE approximately in the middle of the comparison. Despite the increase in model parameters due to the incorporation of the MoE layer, its sparse nature does not significantly escalate the computational load. Figure 12(b) illustrates the correlation between the number of experts and task performance across different modalities. Generally, performance improves with the addition of more experts, plateauing once the count exceeds 16. To achieve a compromise between performance and computational expense, we opted to utilize the top 4 experts out of 16 in our experiments. Figure 12(c) and Figure 13 study the influence of each modality on the top-$k$ chosen experts. For every expert selected, we calculate the number of samples that include a specific modality, weighted by corresponding weight factors from the gating functions. The outcomes are subsequently normalized across modalities. The analysis of Figure 12(c) reveals that predictions across all tasks heavily depend on vital signs and clinical notes. This reliance is attributed to the abundant samples in these two modalities. Despite the notably smaller quantity of CXR, they play more significant roles in the 25-PHE and 48-IHM tasks, which aligns with our findings in Table 5. The results in Figure 13 demonstrate that the modality weight distribution in the MOSI and MOSEI datasets is more “spread out”, with the audio component carrying a greater weight in the MOSEI dataset.
@@ Appendix I Details on Numerical Experiments
We conduct multiple numerical experiments to illustrate the theoretical convergence rates of the MLE $\widehat{G}_{n}$ to the true mixing measure $G_{*}$ under both exact-specified and over-specified settings.
@@ I.1 Experimental Setup
Then, we generate i.i.d samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$ by first sampling $X_{i}$’s from the uniform distribution $\mathrm{Uniform}[0,1]$ and then sampling $Y_{i}$’s from the true conditional density $g_{G_{*}}(Y|X)$ of the Laplace gating Gaussian mixture of experts (MoE) given in equation 4.
Maximum Likelihood Estimation (MLE). A popular approach to determining the MLE $\widehat{G}_{n}$ for each set of samples is to use the EM algorithm Dempster et al. [1977]. However, since there are not any closed-form expressions for updating the gating parameters $\beta_{0i},\beta_{1i}$ in the maximization steps, we have to leverage an EM-based numerical scheme, which was previously used in Nguyen et al. [2024c]. In particular, we utilize a simple coordinate gradient descent algorithm in the maximization steps. Additionally, we select the convergence criterion of $\epsilon=10^{-6}$ and run a maximum of 2000 EM iterations.
@@ I.2 Exact-specified Setting
Under the exact-specified settings, we conduct 5 sample generations for each configuration, across a spectrum of 10 different sample sizes $n$ ranging from $10^{3}$ to $10^{5}$. It can be seen from Figure 3 (left) that the MLE $\widehat{G}_{n}$ empirically converges to the true mixing measure $G_{*}$ under the Voronoi metric $\mathcal{D}_{1}$ at the rate of order ${\mathcal{O}}(n^{-0.49})$, which matches the theoretical parametric convergence rate established in Theorem J.2.
@@ I.3 Over-specified Setting
Under the over-specified settings, we continue to generate 5 samples of size $n$ for each setting, given 10 different choices of sample size $n\in[10^{3},10^{5}]$. From Figure 3 (right), we observe that the MLE $\widehat{G}_{n}$ empirically converges to $G_{*}$ under the Voronoi metric $\mathcal{D}_{2}$ at the rate of order ${\mathcal{O}}(n^{-0.47})$, which aligns with the theoretical parametric convergence rate established in Theorem 3.2.
@@
In this appendix, we study the theoretical behaviors of the MLE under the exact-specified setting, i.e., $k=k_{*}$, of the Laplace gating Gaussian MoE. We demonstrate that under the exact-specified setting, the rate of estimated conditional density function $p_{\widehat{G}_{n}}$ to $p_{G_{*}}$ is parametric $\mathcal{O}n^{-1/2})$ (up to some logarithmic factor).
@@
The proof of Theorem J.1 can be done similarly to that of Theorem 3.1 in Appendix K.2. The result of Theorem J.1 indicates that as long as we can establish the lower bound of the total variation distance between $p_{\widehat{G}_{n}}$ and $p_{G_{*}}$ based on certain loss function between the MLE $\widehat{G}_{n}$ and the true mixing measure $G_{*}$, we directly achieve the rate of the MLE under that loss function.
Voronoi Loss We now define that loss function between the MLE and the true mixing measure for the exact-specified setting:
$\mathcal{D}_{1}(G,G_{*}):=\sum_{j=1}^{k_{*}}\Big{|}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i})-\exp(\beta^{*}_{j})\Big{|}+\sum_{j\in[k_{*}]:|\mathcal{A}_{j}|=1}\sum_{i\in\mathcal{A}_{j}}\exp(\beta_{i})\Phi_{ij}(1,1,1,1).$ (9)
Above, for any $(\rho_{1},\rho_{2},\rho_{3},\rho_{4})\in\mathbb{R}^{4}$, we define $\Phi_{ij}(\rho_{1},\rho_{2},\rho_{3},\rho_{4})=\|W_{i}-W_{j}^{*}\|^{\rho_{1}}+\|a_{i}-a_{j}^{*}\|^{\rho_{2}}+|b_{i}-b_{j}^{*}|^{\rho_{3}}+|\nu_{i}-\nu_{j}^{*}|^{\rho_{4}}$ for any $i\in\mathcal{A}_{j}$ and $j\in[k_{*}]$. We demonstrate in the following theorem that the rate of MLE to the true mixing measure under the Voronoi loss function $\mathcal{D}_{1}$ is $\mathcal{O}(n^{-1/2})$ (up to some logarithmic factor).
@@
Proof of Theorem J.2 is in Appendix K.1. The convergence rate of MLE under the Voronoi loss function $\mathcal{D}_{1}$ implies that the rates of estimating the true parameters $W_{i}^{*},a_{i}^{*},b_{i}^{*},\nu_{i}^{*}$ are also $\mathcal{O}(n^{-1/2})$ (up to logarithmic factors). These rates are comparable to those under the exact-specified setting of softmax gating Gaussian MoE (cf. Theorem 1 in  Nguyen et al. [2023]).